OpenStudy (anonymous):

solve the following equation ( i want with steps) log(x-5) = 1-log(x-14)

OpenStudy (anonymous):

start with \[\log(x-5)+\log(x-14)=1\] then using a property of the log write \[\log((x-5)(x-14))=1\]

OpenStudy (anonymous):

rewrite in equivalent exponential form as \[(x-4)(x-14)=10^1=10\]assuming the log is base ten

OpenStudy (anonymous):

so your final job is to solve the quadratic equation \[(x-4)(x-14)=10\] \[x^2-18x+56=10\] \[x^2-18x+46=0\]etc

OpenStudy (anonymous):

you need to check your answers, because you cannot take the log of negative number

OpenStudy (anonymous):

got it thanks :)))

OpenStudy (anonymous):

yw